A cashier has only five-dollar and twenty-dollar bills in her register. If the register contains 19 bills worth $215, how many five-dollar bills does she have?
Answer by checkley77(12844)

T+F=19 OR T=19-F
20T+5F=215
20(19-F)+5F=215
380-20F+5F=215
-15F=215-380
-15F=-165
F=-165/-15
F=11 FIVE DOLLAR BILLS.
19-11=8 20 DOLLAR BILLS.
PROOF:
20*8+5*11=215
160+55=215
215=215